OpenStudy (hhopke):

Solve for y Options: a.) 8 b.) 16

OpenStudy (hhopke):



OpenStudy (anonymous):

first solve for the triangle on the very left by using pythagorean theorem a^2 + b^2 = c^2

OpenStudy (anonymous):

so it would be (2x2) + b^2 = (4x4)

OpenStudy (hhopke):

I know that the triangles are similar, so we have